# Cultural Transmission, Technology, and Treatment of the Elderly

### v1

@@ Abstract
We discuss the interrelationship between the treatment of the elderly, the nature of production, and the transmission of culture. Respect for the elderly is endogenous. Parents cultivate an interest in consuming culture in their children; when they are older, children compensate their elders proportional to the degree to which their interests were previously cultivated. We show that this model is functionally equivalent to one in which cultural goods are transferred across generations. We focus on the relative well-being of the elderly and use the model to explain patterns in their relative well-being across societies. An important theme is that the cultivation of culture and norms for the respect and support of the elderly bear a nonlinear relationship with many economic variables, such as capital and or land intensity in production. We also discuss the interaction of property rights with production, assets such as productive resources, and relative treatment of the elderly. Insecurity of some types of property rights, such as rights over output, may benefit the elderly, while secure rights over productive resources may also benefit the elderly. We discuss how the elderly could be affected by demographic, technological and policy changes in both developing and developed economies.
@@ 1 Introduction
Aging and the elderly have always been with us. Although this has always and everywhere been true, [^1] A finding of recent anthropological research is that living into old age is common among hunter-gatherers, and was likely common even among early Homo Sapiens (Gurven and Kaplan, 2009). In fact, long lives have been hypothesized to be an evolutionary outcome of the hunter-gatherer lifestyle (Kaplan and Robson, 2002, 2003) there are pronounced differences in the way aging is experienced around the world. In modern and highly developed economies, the elderly are typically able to provide for their own well-being and also receive government-provided benefits. [^2] A United Nations report suggests that, in developed countries, public transfers are the most common means of elderly support in Europe and South America, while in Southeast Asia, the United Kingdom, and the United States, elderly support is mainly financed by asset income. (United Nations Department of Economic and Social Affairs, 2019) In less developed and more traditional societies, including preliterate societies, there is great diversity in the way elderly are cared for and variety in the societal roles the elderly play. The elderly may receive special treatment for the particular knowledge they possess or also for their role as keepers of history and cultural stores. The elderly may benefit from specialized property rights or from the development of property rights in general, as is the case when the elderly gain ownership over productive assets such as land. Sometimes, the elderly may simply take advantage of having lived longer than others, allowing for the accumulation of productive assets and experience.
Demographic and technological changes exert ongoing pressure on the institutions and cultural practices that shape the aging experience. People now live longer and have fewer children than ever before, increasing both the relative and absolute size of the elderly population virtually everywhere. In 1950, there were three people under the age of 5 for every person over the age of 65; projections suggest that by 2050 the opposite will be true (World Health Organization, 2011). Worldwide changes in work, consumption, and savings patterns have consequences for everyone, but these changes have special meaning for the elderly. Some popular writing, such as Diamond (2012, Chapter 6), has drawn attention to aspects of the changing situation of the elderly, and policy makers have begun to craft legislation designed to address the changing situation of the elderly in modern society in some parts of the world. [^3] See https://theweek.com/articles/462230/how-elderly-are-treated-around-world.
Across the development spectrum, and particularly in less developed economies, the elderly play a critical role in the maintenance and transmission of culture. Therefore, it is natural to ask the following. What implications might the changes that profoundly impact the elderly have for cultural transmission and retention? [^4] As one report on global aging puts it, “…some countries may grow old before they grow rich.” (World Health Organization, 2011, p. 5) Much of the growth of the elderly population in the coming decades is expected to occur in less developed countries, where the condition of the elderly is closely linked with tradition and culture. [^5] World Health Organization (2011) estimates that the world’s population 65 or older will grow from 524 million in 2010 to 1.5 billion in 2050, most of this growth occurring in developing countries. United Nations Department of Economic and Social Affairs (2019) estimates that by 2050, 1 in 6 people in the world will be over the age of 65.
We construct a simple model that provides insight into the forces driving the diversity of elderly treatment across societies and across the spectrum of development, from hunter-gatherer societies to modern economies. Since the model sheds light on some of the aspects of material life that have led to cultural practices that support the elderly, it provides insight into how technological and demographic change impact the way elderly are treated. At the same time, the model provides insight into how elderly treatment and cultural transmission are intertwined and therefore offers an understanding of how transmission of culture itself interacts with technological and demographic change.
@@ 2 Literature
Formal analysis of the condition of the elderly has attracted the attention of economists at least since mathematical models of growth became an integral part of economics. Aging, savings behavior, and capital accumulation are intimately connected in many, if not most, of the core dynamic economic models. The life cycle model of Modigliani (1966) is explicitly built on the relationship between aging, savings behavior, and accumulation. The overlapping generations (OLG) model developed in Samuelson (1958) and Diamond (1965) also intimately binds savings, demographics, and accumulation.
The OLG model has proven to be useful in analyzing the interrelationship between capital accumulation, aging, and family structure, and has also been used to study the role that inter vivos transfers play in an economy, forging a direct link between generations. [^6] A review of literature on altruism and inter-generational transfers is Michel et al. (2006). Boldrin and Jones (2002) and Nishimura and Zhang (1995) reverse the usual parent-child altruism of Barro and Becker (1989), and instead study situations in which children are altruistic towards their parents. These models inform elderly treatment; in an environment with child-to-parent altruism, transfers to the elderly emerge naturally, as do investments in fertility and children’s education, which may be altruistic but may also result from self-interested behavior on the part of the elderly, as such investments may enhance subsequent transfers to the elderly.
Some research, following models of strategic interfamilial behavior like Bernheim et al. (1985), posits that intergenerational transfers to and from the elderly may play a critical strategic role in the maintenance of the family fabric. Cigno (1993) show how transfers from the current working-age population to both young and old may be part of a self-enforcing set of family rules; a chief finding of Cigno (1993) is that the existence of such rules is sensitive to the existence of capital markets and social support systems. [^7] Kaplan and Robson (2003) have suggested that an important aspect of human evolution might require such intergenerational capital markets. As Laferrére (1999) documents in her survey of models of intergenerational transmission, there are a wide variety of models in which strategic motives both in transferring wealth between family members of different generations and actively encouraging altruistic behavior between family members emerge. A prominent example is Morand (1999), who discusses the relationship between fertility and growth in a model in which the elderly make fertility decisions with an eye on the support of the elderly, linking the literature on demographic transition (Galor and Weil, 2000) and the support of the elderly. Blackburn and Cipriani (2006) develop a model in which societies transition from economies characterized by high fertility rates, low levels of development, and intergenerational transfers that flow from young to old, to economies with low fertility rates and high levels of development, with transfers flowing from old to young.
A literature focusing on cultural transmission of familial norms through parental socialization efforts and the presence of coincident traits in the population, following the approach introduced in Bisin and Verdier (2000, 2001), has also been applied to shed light on the evolution of elderly treatment. Olivera (2013) studies a model in which agents have a genetic disposition to care for the elderly (or not), and shows how support for old age relates to fertility decisions. In this type of model, norms of elderly support may persist, even as fertility declines and life expectancy increases. Varvarigos (2021) explores a model similar in spirit to Olivera (2013) in which adults influence children’s preferences to be like their own, but with a focus on the implications for development. Varvarigos’s model produces equilibria in which low levels of economic development, strong family ties, and intergenerational transfers coexist; the model is also capable of producing equilibria in which intergenerational transfers are not important, family bonds are not as strong, and the economy functions at higher levels of development. Cigno et al. (2017) and Cigno et al. (2020) combine marital classification and trait transmission with aspects of intergenerational self-enforcing contracts originally developed in Cigno (1993). This work shows how changes in education levels and wages can reduce the prevalence of family contracts focusing on the elderly.
The Bisin and Verdier (2000, 2001) model style has also been used to study the relationship between social support systems and elderly treatment. Cremer et al. (2013), for example, presents a model in which parents spend time with children as a means to gain support for a possible future disability and study the interaction of support with policy. Ponthiere (2013) discusses a replacement effect in which state-provided long-term care can modify the transmission of values and affect family altruism.
Although our model is similar in spirit to these models, the basic mechanism underlying culture transmission and how it relates to support for the elderly is different. In our model, middle-aged agents work to cultivate an interest in their children in the consumption of a cultural good provided by the elderly, as is reminiscent of Cozzi (1998). Adopting this approach firmly focuses our paper on the role in which cultural transmission is related to other savings decisions that the agents might make, rather than working through a secondary channel in which agents care about the propagation of their genetic traits. It also forms a direct link between the appreciation and maintenance of culture and the treatment of the elderly. Under some circumstances, our approach is then isomorphic to an approach in which agents invest in cultivating an interest in culture as a means of storing wealth for consumption in the future. Since current middle-aged will be elderly in the future, if they cultivate interest in cultural consumption among their children in the present, future consumers of their cultural goods may in part sustain them in the future.
There has been extensive empirical work on altruism toward the elderly. Early work on intergenerational transfers within the family found that forming hard conclusions about the motivations driving transfers is elusive (Laferrére, 1999), as purely altruistic motives for giving are difficult to separate from self-interested motives. In a recent example, Klimaviciute et al. (2017) apply data from a large European dataset on children’s and parents’ behavior concerning intergenerational transfers to test models of altruism towards the elderly, to determine how support policies could interact with the interfamilial provision of elderly care. Their findings support “moderate altruism” as the driving force behind inter vivos transfers to the elderly.
Therefore, there have been a variety of different approaches to studying the evolving nature of elderly treatment and how elderly treatment interacts with strategic, altruistic and technological aspects of the economy. Our approach extends these approaches in a way that melds strategic and altruistic aspects of elderly treatment, and that also explicitly links economic growth, the development of property rights, and the nature of production to the treatment of the elderly. Most previous research [^8] But not all. Varvarigos (2021), for example, is a general equilibrium analysis. relies on a partial equilibrium approach and does not seek to explore deeper trends in elderly treatment, production technology and economic growth in a way that incorporates a typical growth model such as that espoused in Acemoglu (2009, Chapter 9).
Furthermore, the existing literature does not explicitly focus the analysis on the treatment of the elderly. The cross-cultural ethnographic literature points out several critical factors of the economy driving the treatment of the elderly, such as the origins and development of property rights. Our approach allows us to build a model capable of replicating this evidence, while along the way developing working measures of elusive concepts such as “elderly treatment” and “elderly prestige.” Finally, our model tightly links technical change and elderly treatment and allows us to see how preferences, elderly treatment, and technology co-evolve.
@@ 3.1 Cross-cultural research
The logical place to begin a discussion of cross-societal analysis of the treatment of the elderly is the seminal work of Simmons (1970) (originally written in 1945), who surveyed the treatment of the elderly in a sample of 71 different societies. Simmons’ sample included hunter-gatherers, pastoral and nomadic herding peoples, and subsistence agriculturalists from all over the world. In constructing these data, Simmons aimed to create a sample of societies that were at once technologically, geographically, and environmentally diverse. [^9] Simmons’s approach mirrors that employed by G. P. Murdock and his co-researchers in the construction of the Ethnographic Atlas (Murdock, 1969), and Murdock and White in their construction of the Standard Cross-Cultural Sample (Murdock and White, 1969), and also resembles the data used in Pryor (1973). See Baker and Jacobsen (2007b), Baker and Jacobsen (2007a), or Baker and Miceli (2005) for an application and further description of cross-cultural data within the context of economic modeling. This sample allowed Simmons to develop a fairly comprehensive picture of regularities in elderly treatment across societies and thus form a picture of how the role of the elderly changes along with the nature of subsistence and societal complexity. In addition to setting the stage for subsequent work on cross-cultural treatment of the elderly, Simmons’s survey also provides a useful snapshot of elderly treatment around the world before the onset of global modernization. The following subsections organize and explain some of the key results of this exercise. Our summary of this information and the organization of our summary also relies on the review of Simmons’ work in Holmes (1976) and the discussion of Arnhoff et al. (1964).
@@ 3.1.1 Food Sharing and Taboos
In societies without well-developed property rights (perhaps subsistence hunter-gatherer societies), practices such as communal food sharing and food taboos provide consumption security for the elderly. Elderly consumption becomes more secure as property rights in land and in other property, such as livestock, develop.
@@
Most cultures earning subsistence by hunting and gathering also usually have some system for food sharing. The existence of sharing has been attributed to a wide variety of motives - insurance, exchange, even resource management - making it difficult to surmise “the” reason output is shared. [^10] See Baker and Swope (2021) for an overview of the many theories of sharing among hunter-gatherers. Kelly (2013, Chapter 6) has an extensive discussion of sharing across hunter-gatherer populations. Nonetheless, a common feature of food sharing is that the elderly tend to be included, so food sharing provides food for the elderly. Thus, to use the terminology of Lee (2003), sharing redistributes “vertically” from adults to both the elderly and children, as well as “horizontally” - that is, between hunters. It is also true that many hunter-gatherers share access to resources, so it seems that the lack of individual ownership rights over the food output that is produced is correlated with the lack of ownership over land and other resources. [^11] This of course does not mean hunter-gatherers aren’t territorial at the group level; see (Baker, 2003).
The practice of tabooing food in a way advantageous to the elderly, referred to as “Victual Deference” by Silverman and Maxwell (1978), is common among hunter-gatherers living in a wide variety of environments. Simmons describes taboos as “…a social expedient of particular advantage to the aged…” and also notes that “…while it cannot be demonstrated that these oldsters invented such social expedients in behalf of their own palates, many of the tabooed foods were regarded as quite tasty, and frequently the aged have been observed manipulating such taboos to their personal advantage.”(Simmons, 1970, p. 26) Food taboos are a useful illustration of a core principle embodied in the last item on the list, as they are very likely ingrained in individuals through conscious effort. They therefore represent a telling example of a cultural practice that is taught and transmitted and, at the same time, may actually serve no other practical purpose than to secure subsistence for the elderly. Simmons gives food tabooing examples from a diverse group of societies around the world, including Native American groups such as the Greenland Inuit, Omaha, and Pomo; the Iban of Borneo; and the Arunta from the Australian desert, to name a few.
Examples beyond Simmons’s work of taboos are easy to find. Food taboos feature prominently among the Ju/’hoansi, a hunting and gathering people that inhabit the Kalahari desert in Namibia and Botswana, who have been the subject of a wealth of research. [^12] In English, this would be something like ‘juntwasi.’(Rosenberg, 2009) In earlier writings, they were known as the !Kung. See, e.g., Lee (1979). For example, “elders are felt to have special powers that permit them to eat certain foods (e.g., ostrich eggs) considered too dangerous for younger people to consume.” (Rosenberg, 2009, p. 33). McArdle and Yeracaris (1981) provide another example from the Chukchee people of Siberia that they attribute to Bogoras (1904), in which it is believed that Reindeer milk is harmful to young people. An interesting variation on this theme comes from the Pukapukan people of the Cook Islands, where there exists a “culturally conditioned system of both food sharing and food taboos” (Simmons, 1970, p. 10) that favored the elderly, in which the elderly controlled the distribution of food and consequently were able to direct to themselves some of the most nutritious animal parts, such as belly fats and internal organs.
@@ 3.1.2 Property rights
While in the case of food sharing, it might be said that ill-defined property rights benefit the elderly, Simmons observes that “property rights have been lifesavers for the aged,” (Simmons, 1970, p. 36), implying that the definition of other types of property rights benefit the elderly. Simmons overlaps a few different concepts of property rights, and the distinction is important. As discussed in the previous section, nonexclusive rights over food help the elderly, but it is also the case that ownership over assets, as it emerges, may benefit the elderly.
Simmons finds that among people who rely on agriculture or animal husbandry (or both) for subsistence, the elderly are also afforded respect as among hunter-gatherers, but for seemingly different reasons, as in these cases respect likely owes to accumulation of wealth. Simmons documents the accumulation of assets such as working animals and livestock among the elderly in societies such as Navajo (Southwestern North American Desert), Chukchi (Siberia near the Bering Strait), Yakut (Russian far east), and Akamba (Kenya), to name a few. Peoples who rely primarily on agriculture coupled with land ownership also seem to feature respect for the elderly coincident with accumulation of assets. As (Simmons, 1970, p. 44) indicates, “Firmly entrenched property rights which enhanced the powers of the aged have also been very common among all cultivators of the soil.”

### v2

# Cultural Transmission, Property Rights, and Treatment of the Elderly

@@ Abstract
We examine how production and the development of property rights interact with cultural transmission to shape the treatment of the elderly across societies. Our model posits that respect for the elderly arises endogenously: parents invest in cultivating cultural values in their children, who later reciprocate in proportion to this investment. We show that this model is functionally equivalent to one in which cultural goods are transferred by the elderly. We focus on the distinct roles of property rights, finding that while insecure output rights may promote elderly welfare, secure rights over productive resources can have comparable benefits. The model reveals a nonlinear relationship between cultural sophistication, property rights, and economic factors such as the capital and land intensity of production, driving variations in elderly well-being across societies. Finally, we consider how the model suggests demographic, technological, and policy changes influence elderly well-being across the spectrum of development.
@@ 1 Introduction
Aging and the elderly have always been part of human existence. Living into old age is common among hunter-gatherers and was likely even common among early Homo sapiens, suggesting that long lifespans may be an evolutionary adaptation to the hunter-gatherer lifestyle (Gurven and Kaplan, 2009; Kaplan and Robson, 2002, 2003). However, the experience of aging varies significantly across societies. In modern economies, the elderly often sustain themselves through personal means and government-provided benefits (United Nations Department of Economic and Social Affairs, 2019). By contrast, in traditional societies, including preliterate ones, the elderly’s roles and treatment are diverse. They may be revered for their knowledge, act as cultural custodians, or benefit from specialized or evolving property rights. In some cases, their longevity enables them to accumulate productive assets and experience.
Demographic and technological shifts are reshaping institutions and cultural practices surrounding aging. Increased longevity and declining fertility have dramatically expanded the elderly population worldwide, both in total and relative to younger ages. The global population aged 65 or older is expected to rise from 524 million in 2010 to 1.5 billion by 2050 (World Health Organization, 2011). In 1950, there were three children under five for every person over 65; by 2050, this ratio is projected to reverse (World Health Organization, 2011). These changes influence work, consumption, and savings patterns, particularly for the elderly. Popular works, such as Diamond (2012, Chapter 6), highlight aspects of this transformation, and some governments are responding with targeted policies (United Nations Department of Economic and Social Affairs, 2019). [^1] See https://theweek.com/articles/462230/how-elderly-are-treated-around-world.
Across all economic development levels, but especially in less developed economies, the elderly play a key role in preserving and transmitting culture. This raises critical questions: how might demographic and technological changes affecting the elderly influence cultural transmission? (World Health Organization, 2011, p. 5). Notably, much of the anticipated growth in the elderly population will occur in lower income countries, where tradition and culture strongly shape their condition (World Health Organization, 2011; United Nations Department of Economic and Social Affairs, 2019).
We present a simple model to examine the factors shaping elderly treatment across societies, from hunter-gatherers to modern economies. The model shows how material conditions shape cultural practices supporting the elderly and how technological and demographic shifts transform these practices. It highlights the interaction between property rights over productive assets and output, elderly treatment, and cultural transmission. A key result is the emergence of what anthropologists refer to as a ‘curvilinear’ pattern: societies midway between traditional and modern economies are predicted to have the strongest cultural traditions of reverence for the elderly.
The rest of this paper is organized as follows. In Section 2, we relevant models from the literature on the economics of aging, and how they relate to property rights, cultural transmission, and modernization. In Section 3, we discuss aspects of the cross-cultural literature on aging and treatment of the elderly. In Section 4, we present our model, which emphasizes property rights, and shows how a ‘curvilinear’ pattern in cultural concern for the elderly emerges. Section 5 discusses implications of results and concludes.
@@ 2 Literature
Economic analysis of aging has long focused on the connection between aging, savings behavior, and capital accumulation, central to many dynamic economic models. For example, the life-cycle model of Modigliani (1966) and the overlapping generations (OLG) models of Samuelson (1958) and Diamond (1965) explicitly link these factors. The OLG model has proven particularly useful for studying how aging interacts with family structure, intergenerational transfers, and capital accumulation (Michel et al., 2006). Some models, such as those by Boldrin and Jones (2002) and Nishimura and Zhang (1995), invert the traditional altruism paradigm, exploring child-to-parent altruism, which naturally generates elderly support and investment in children’s education and fertility.
Other research emphasizes the strategic role of intergenerational transfers in family cohesion. For example, Cigno (1993) shows that family-based transfers can be self-enforcing under specific economic conditions, such as the availability of capital markets and social support systems (Kaplan and Robson, 2003). Studies like Morand (1999) and Blackburn and Cipriani (2006) further explore the connection between fertility, growth, and elderly support, highlighting transitions from high-fertility, traditional economies to low-fertility, modern ones, with changing patterns of intergenerational transfers.
Models of cultural transmission, such as those by Bisin and Verdier (2000, 2001), offer additional insights into elderly treatment. Olivera (2013) demonstrates how norms of elderly care can persist despite declining fertility and rising life expectancy. Similarly, Varvarigos (2021) links strong family ties, intergenerational transfers, and traditional means of production to evolving societal norms. Cremer et al. (2013) and Ponthiere (2013) analyze the interaction between social support systems and familial altruism, showing how state-provided care can influence family dynamics and value transmission.
Our approach builds on these foundations but focuses on how cultural transmission interacts with elderly treatment. Unlike models centered on genetic or altruistic motives, we examine how middle-aged individuals cultivate their children’s interest in cultural goods provided by the elderly, creating a direct link between cultural preservation and elderly support. This perspective aligns with Cozzi (1998), where cultural investment serves as a form of wealth storage for future consumption.
Empirical work on altruism toward the elderly supports the complexity of these dynamics. While early studies struggled to disentangle altruistic and self-interested motives (Laferrére, 1999), recent research, such as Klimaviciute et al. (2017), finds evidence of ‘moderate altruism’ driving intergenerational transfers.
Our model advances this literature by integrating strategic and altruistic aspects of elderly treatment with economic growth, property rights, and production. Unlike most prior research, which uses partial equilibrium approaches, we adopt a general equilibrium framework, as in Acemoglu (2009, Chapter 9). This approach allows us to explore how preferences, technology, and elderly treatment co-evolve and to develop measures of ‘elderly treatment’ and ‘elderly prestige.’ It also connects technological change with the evolving status and treatment of the elderly, providing a comprehensive framework for understanding these dynamics across cultures and economies.
@@ 3.1 Cross-cultural research
A foundational study in cross-societal analysis of elderly treatment is Simmons (1970) (originally published in 1945), which surveyed 71 societies, including hunter-gatherers, pastoralists, and subsistence agriculturalists from diverse technological, geographical, and environmental conditions. Simmons aimed to capture a broad cross-section of societies to identify patterns in elderly treatment and how these patterns vary with subsistence strategies and societal complexity. [^2] Simmons’s methodology resembles the approaches used by G. P. Murdock and collaborators in the Ethnographic Atlas (Murdock, 1969) and the Standard Cross-Cultural Sample (Murdock and White, 1969), as well as data utilized in Pryor (1973). For applications of cross-cultural data in economic modeling, see Baker and Jacobsen (2007b, a); Baker and Miceli (2005).
Simmons’s work provides a comprehensive snapshot of elderly treatment across societies before the influence of global modernization and established a framework for subsequent cross-cultural studies. It highlights how the role of the elderly evolves with societal complexity and subsistence strategies. Our summary draws on Simmons’ findings, the review by Holmes (1976), and the discussion in Arnhoff et al. (1964).
@@ 3.1.1 Food Sharing and Taboos
In societies without well-developed property rights, such as subsistence hunter-gatherer groups, communal food sharing and food taboos provide consumption security for the elderly. As property rights in land and livestock develop, elderly consumption becomes more secure (Simmons, 1970, Ch. 2).
@@
Food sharing is common among hunter-gatherers, motivated by factors such as insurance, exchange, or resource management, though pinpointing a singular reason is difficult. [^3] See Baker and Swope (2021) for an overview of sharing theories among hunter-gatherers, and Kelly (2013, Chapter 6) for an extensive discussion. Sharing typically redistributes resources ‘vertically’ from adults to both the elderly and children, and ‘horizontally’ among hunters (Lee, 2003). Moreover, hunter-gatherers often lack individual ownership rights over food and land, correlating shared food output with communal access to resources. [^4] However, hunter-gatherers can still be territorial at the group level; see (Baker, 2003).
Food taboos advantageous to the elderly, described as ‘Victual Deference’ by Silverman and Maxwell (1978), are widespread among hunter-gatherers in various environments. Simmons noted that these taboos serve as “a social expedient of particular advantage to the aged” and are often manipulated by the elderly for personal benefit (Simmons, 1970, p. 26). Examples from Simmons include the Inuit, Omaha, Pomo, Iban, and Arunta, where tabooed foods often favored the elderly. Additional examples include the Ju/’hoansi of the Kalahari Desert, where elders are believed to have powers enabling them to consume foods considered too dangerous for younger individuals, such as ostrich eggs (Rosenberg, 2009, p. 33). Similarly, among the Chukchee of Siberia, reindeer milk is deemed harmful to the young (McArdle and Yeracaris, 1981; Bogoras, 1904). In the Cook Islands, the Pukapukan people maintained a system of food sharing and taboos that allowed elders to control distribution and reserve nutritious parts, such as belly fats and internal organs, for themselves (Simmons, 1970, p. 10).
@@ 3.1.2 Property rights
While nonexclusive food sharing benefits the elderly, Simmons observes that “property rights have been lifesavers for the aged” (Simmons, 1970, p. 36), suggesting that well-defined property rights over assets also benefit them. Simmons distinguishes between these types of rights, emphasizing that while nonexclusive rights over food help sustain the elderly, ownership of assets like land or livestock provides additional advantages.
Simmons documents asset accumulation by the elderly, such as working animals and livestock, in groups like the Navajo (Southwestern North America), Chukchi (Siberia), Yakut (Russian Far East), and Akamba (Kenya). In agricultural societies with land ownership, respect for the elderly often aligns with asset accumulation. As Simmons notes, “Firmly entrenched property rights which enhanced the powers of the aged have also been very common among all cultivators of the soil” (Simmons, 1970, p. 44).